OpenStudy (anonymous):

on monday, Roger drove to work in 20 minutes. on tuesday, he averaged 15 miles per hour more, and it took him 6 minutes less to get to work. how far in miles does he travel to work?

OpenStudy (anonymous):

x = distance y = velocity x/y = 20/60 x/y+15 = 14/60 x = 11.7 miles y = 35 mph

OpenStudy (welshfella):

first we need to convert minutes to hours as we are dealing with mph. 20 minutes = 1/3 hour and 14 minutes = 7/30 hours v = speed , x = distance v = s/t we have the 2 equations v = x / 1/3 = 3x v + 15 = x / (7/30) = 30x/7 substitute v = 3x in 2nd equation:- 3x + 15 = 30x/7 21x + 105 = 30x 9x = 105 x = 11.7 miles

OpenStudy (welshfella):

* thats v = x/t